OpenStudy (theraggedydoctor):

\(x^2-2x+3=x^3+2x+x^2\) is equivalent to \(0\) \(2x^2-4x=0\) \(-x^3+4x-3=0\) \(x^3-2x^2-3=0\) \(x^3+4x-3=0\)

OpenStudy (loser66):

YOur idea?

OpenStudy (welshfella):

subtract x^2 , -2x and 3 in turn from the left hand side and see what you get

OpenStudy (theraggedydoctor):

Okay. Hold on..

OpenStudy (mayankdevnani):

what do you think ?

OpenStudy (theraggedydoctor):

I really dont know.

OpenStudy (mayankdevnani):

\[\large \bf x^2-2x+3=x^3+2x+x^2\] first subtract x^2 to both sides,

OpenStudy (mayankdevnani):

what we get ?

OpenStudy (theraggedydoctor):

\(x^2\) cancels out..

OpenStudy (anonymous):

Equivalent to 0

OpenStudy (theraggedydoctor):

How? Show me..

OpenStudy (mayankdevnani):

when x^2 cancels out , we get \[\large \bf -2x+3=x^3+2x\] Add 2x to both the sides

OpenStudy (mayankdevnani):

what we get ?

OpenStudy (anonymous):

Use, http://mathpapa.com/algebra-calculator.html

OpenStudy (theraggedydoctor):

We get.. \(-2x+2x+3=x^3\)

OpenStudy (mayankdevnani):

nope ! try again

OpenStudy (theraggedydoctor):

But I added the \(2x\) from the other side..

OpenStudy (theraggedydoctor):

Oh.. to both sides. @mayankdevnani

OpenStudy (theraggedydoctor):

I got \(4x\) on the right side.

OpenStudy (mayankdevnani):

so we get \[\large \bf 3=x^3+4x\] now subtract 3 to both sides,we get \[\large \bf x^3+4x-3=0=\color{red}{Answer}\]

OpenStudy (mayankdevnani):

hope you understand ! @TheRaggedyDoctor

OpenStudy (theraggedydoctor):

Oh, okay. Thank you. Yes! I understood.

OpenStudy (mayankdevnani):

your welcome :)